### v1

@@ A-E Proof of Optimal Stopping Theorem 4
$\displaystyle V(\pi)=\min\{0,\chi(y)-g+\rho\mathds{E}V(\pi)\}$
Further, Lemma 2 of [7] shows that the incentive function is decreasing. From the definition of First order stochastic dominance and Proposition 1 of [7] $\mathds{E}V_{m}(\pi)$ is decreasing in $\pi$. Therefore $V_{m}(\pi)$ is decreasing which implies $V(\pi)$ is decreasing. Let $V(0)$ and $V(1)$ be the value at $\pi=[1,0]$ and $\pi=[0,1]$ respectively which makes $\mathds{E}V(0)=V(0)$ and $\mathds{E}V(1)=V(1)$.
@@ A-F Proof of Submartingale Result Theorem 5
$\hat{\nu}(\pi)=\begin{cases}\chi(\pi)-\epsilon&\text{if}\ \pi(2)\leq\bar{\pi}^{*}(2)\\ \chi(\pi)&\text{if}\ \pi(2)>\bar{\pi}^{*}(2)\\ \end{cases}$
@@
$\displaystyle H_{k}=\begin{cases}0\text{if}\ \pi_{k-1}(2)&\leq\bar{\pi}^{*}(2)\\ 1\text{if}\ \pi_{k-1}(2)&\geq\bar{\pi}^{*}(2)\end{cases}$
@@
${\mathds{P}\left[\max_{1\leq k\leq T}p_{k}\geq C\right]\leq{\frac{\mathds{E}[{\max}(p_{T},0)]}{C}}}={\frac{\mathds{E}[p_{T}]}{C}}$
@@ B-B4 FNSPID
$\frac{\mathsf{C}_{k+1}}{\mathsf{C}_{k}}-\frac{\mathsf{C}_{k}}{\mathsf{C}_{k-1}},$
where $\mathsf{C}_{k}$ is the close price at time $k$. The close price is the stock price at the end of the trading day.
@@ B-D2 Likelihood Neural Network (Restricted Boltzmann Machine)
[Uncaptioned image] Adit Jain received his Bachelor of Technology in Electronics and Communication Engineering from the Indian Institute of Technology, Guwahati in 2022 where he received the Institute Silver Medal. He is currently a graduate student at Cornell University. He has internship experience at Adobe Research and Goldman Sachs. His research interests include structural results for episodic reinforcement learning, online learning, and high-dimensional linear bandits. His research is focused on applications in federated learning, large language model agents, and active learning. He is a recipient of the Data Science Fellowship by the Cornell Center for Social Sciences.

### v2

@@ A-E Proof of Optimal Stopping Theorem 4
$\displaystyle V(\pi)=\min\{0,\chi(y)-g+\rho\mathds{E}V(\pi)\}.$
Further, Lemma 2 of [2] shows that the incentive function is decreasing. Denote $V_{m}$ to be the $m-$th iterate of the value iteration algorithm [5] which iteratively converges to the value function $V(\pi)$. The iterates are given by,
$\displaystyle V_{m+1}=\min\{0,\chi(y)-g+\rho\mathds{E}V(\pi)\}.$
From the definition of first order stochastic dominance and Proposition 1 of [2] $\mathds{E}V_{m}(\pi)$ is decreasing in $\pi$.
Therefore $V_{m}(\pi)$ is decreasing which implies $V(\pi)$ is decreasing. Let $V(0)$ and $V(1)$ be the value at $\pi=[1,0]$ and $\pi=[0,1]$ respectively which makes $\mathds{E}V(0)=V(0)$ and $\mathds{E}V(1)=V(1)$.
@@ A-F Proof of Submartingale Result Theorem 5
$\hat{\nu}(\pi)=\begin{cases}\chi(\pi)-\epsilon&\text{if}\ \pi(2)\leq\bar{\pi}^{*}(2)\\ \chi(\pi)&\text{if}\ \pi(2)>\bar{\pi}^{*}(2)\\ \end{cases}.$
@@
$\displaystyle H_{k}=\begin{cases}0\ \text{if}\ \pi_{k-1}(2)&\leq\bar{\pi}^{*}(2)\\ 1\ \text{if}\ \pi_{k-1}(2)&\geq\bar{\pi}^{*}(2)\end{cases}.$
@@
${\mathds{P}\left[\max_{1\leq k\leq T}p_{k}\geq C\right]\leq{\frac{\mathds{E}[{\max}(p_{T},0)]}{C}}}={\frac{\mathds{E}[p_{T}]}{C}}.$
@@ B-B4 FNSPID
$\frac{C_{k+1}}{C_{k}}-\frac{C_{k}}{C_{k-1}},$
where $C_{k}$ is the close price at time $k$. The close price is the stock price at the end of the trading day.
@@ B-D2 Likelihood Neural Network (Restricted Boltzmann Machine)
[Uncaptioned image] Adit Jain received his Bachelor of Technology in Electronics and Communication Engineering from the Indian Institute of Technology, Guwahati in 2022 where he received the Institute Silver Medal. He is currently a graduate student at Cornell University. He has internship experience at Adobe Research and Goldman Sachs. His research interests include structural results for episodic reinforcement learning, online learning, and high-dimensional linear bandits. His research is focused on applications in federated learning, large language model agents, and active learning. He is a recipient of the Data Science Fellowship graciously offered by the Cornell Center for Social Sciences.